# Lattice point enumeration of polytopes associated to integer compositions

### v2

@@ Abstract.
An $n$-dimensional lattice polytope ${\mathcal{Q}}_{\sigma}$ can be associated to any composition $\sigma$ of a positive integer $n$, as a special case of constructions due to Pitman–Stanley and Chapoton. The entries of the $h$-vector of $\sigma$, introduced by Chapoton, enumerate the lattice points in ${\mathcal{Q}}_{\sigma}$ by the number of their zero coordinates. Chapoton conjectured that this vector is equal to the $h$-vector of a simplicial polytope. This paper proves that the $h$-vector of $\sigma$ is gamma-positive, hence palindromic and unimodal, and confirms certain other conjectures of Chapoton on the lattice point enumeration of composition polytopes. A combinatorial interpretation of their $h^{\ast}$-polynomials is deduced.
@@ 1. Introduction
The face lattice of composition polytopes is well understood. As observed in [12, Theorem 20], ${\mathcal{Q}}_{\sigma}$ is combinatorially isomorphic to a product of simplices, the dimensions of which are the parts of $\sigma$. The lattice point enumeration of composition polytopes, which is more interesting, has been studied for the more general classes of arbor polytopes [5] and Pitman–Stanley polytopes [12]. Chapoton defined the $h$-vector of a composition $\sigma$ of $n$ (more generally, of an arbor) as the sequence $h(\sigma)=(h_{i}(\sigma))_{0\leq i\leq n}$, where $h_{i}(\sigma)$ is the number of points in ${\mathcal{Q}}_{\sigma}\cap{\mathbb{N}}^{n}$ having exactly $i$ nonzero coordinates (this vector should not be confused with the $h$-vector of the simple polytope ${\mathcal{Q}}_{\sigma}$). We note that $h_{0}(\sigma)=h_{n}(\sigma)=1$. The following intriguing conjecture has been stated more generally for the class of arbor polytopes [5, Conjecture 0.1]; we refer to the discussion in [5, Section 0.1] for the motivation behind it.
@@
In particular, $h(\sigma)$ should be palindromic and unimodal, meaning that $h_{i}(\sigma)=h_{n-i}(\sigma)$ for all $0\leq i\leq n$ and $h_{0}(\sigma)\leq h_{1}(\sigma)\leq\cdots\leq h_{\lfloor n/2\rfloor}(\sigma)$. Moreover, it is suggested in [5, Section 0.1] that $h(\sigma)$ should in fact be equal to the $h$-vector of an $n$-dimensional flag simplicial polytope. Thus, in view of Gal’s conjecture [6, Conjecture 2.1.7] (see also [1, Section 3]) on gamma-positivity, it is reasonable to expect that the generating polynomial $h(\sigma,t):=\sum_{i=0}^{n}h_{i}(\sigma)t^{i}$ of $h(\sigma)$ is $\gamma$-positive, meaning that
@@
for some nonnegative integers $\gamma_{i}(\sigma)$. The first main result of this paper shows that this is indeed the case.
@@
The polynomial $h(\sigma,t)$ is $\gamma$-positive for every composition $\sigma$. In particular, the vector $h(\sigma)$ is palindromic and unimodal.
Our second main result confirms another curious conjecture of Chapoton [5], concerning the Ehrhart polynomial of composition polytopes. We recall (see, for instance, [10, Section 4.6]) that the Ehrhart polynomial ${\rm Ehr}(Q,t)$ of a lattice polytope ${\mathcal{Q}}\subset{\mathbb{R}}_{\geq 0}^{n}$ has the defining property that ${\rm Ehr}({\mathcal{Q}},m)=\#(m{\mathcal{Q}}\cap{\mathbb{N}}^{n})$ for every $m\in{\mathbb{N}}$. The zeta polynomial (see [10, Section 3.12]) ${\mathcal{Z}}(P,t)$ of a finite poset $(P,\preceq)$ has the defining property that ${\mathcal{Z}}(P,m+1)$ is equal to the number of $m$-element multichains in $P$, meaning sequences $(p_{1},p_{2},\dots,p_{m})$ of elements of $P$ such that $p_{1}\preceq p_{2}\preceq\cdots\preceq p_{m}$, for every $m\in{\mathbb{N}}$. Given a composition $\sigma$ of $n$, Chapoton [5, Section 1] has considered the set $P_{\sigma}$ of lattice points in ${\mathcal{Q}}_{\sigma}$ as a subposet of the product poset ${\mathbb{N}}^{n}$.
@@
([5, Conjecture 3.1]) We have ${\rm Ehr}({\mathcal{Q}}_{\sigma},t)={\mathcal{Z}}(P_{{\rm rev}(\sigma)},t+1)$ for every composition $\sigma$, where ${\rm rev}(\sigma)$ is the composition obtained from $\sigma$ by reversing the order in which the parts of $\sigma$ appear.
Here is another intriguing conjecture from [5], stated there for all arbor polytopes (see [5, Conjecture 2.2]), on which no progress is made in this paper.
@@
Theorem 1.2 is proven in Section 2; the proof yields an explicit combinatorial interpretation of the numbers $\gamma_{i}(\sigma)$ which appear in Equation (2) (see Proposition 2.2). Section 2 also shows that $h(\sigma)=h({\rm rev}(\sigma))$ for every composition $\sigma$. Two proofs of Theorem 1.3, one essentially bijective and another based on the results of [12, Section 3], are given in Section 3 and a combinatorial interpretation of the $h^{\ast}$-polynomial of ${\mathcal{Q}}_{\sigma}$ is deduced via the theory of pure shellability of posets (see Proposition 3.2). While this paper is restricted to composition polytopes, we do hope that some of the arguments presented here can be extended to more general classes of lattice polytopes.
Throughout this paper, we set ${\mathbb{N}}=\{0,1,2,\dots\}$ and $n]=\{1,2,\dots,n\}$ for $n\in{\mathbb{N}}$. We denote by #S\#S the cardinality of a finite set $S$ and write $\sigma\models n$ when $\sigma$ is a composition of $n$. Our notation and terminology about posets follows [10, Chapter 3]. Background on Ehrhart theory and face enumeration of simplicial complexes and polytopes can be found in [3, 7] [10, Chapter 4] and [9], respectively.
@@ 2. Gamma-positivity
This section proves Theorem 1.2 and shows (see Proposition 2.5) that $h(\sigma)=h({\rm rev}(\sigma))$ for every composition $\sigma$.
@@
(see [1, Theorem 2.32] for a generalization in the setting of finite Coxeter groups). ∎
For the remainder of this section we represent elements of ${\mathbb{N}}^{n}$ with sum of coordinates at most $n$ by lattice paths in the $xy$-plane as follows. A point $(a_{1},a_{2},\dots,a_{n})\in{\mathbb{N}}^{n}$ with $a_{1}+a_{2}+\cdots+a_{n}\leq n$ is represented by the lattice path with initial vertex (1,0)(-1,0), terminal vertex $(n,n+1)$ and east (E) or north (N) unit steps, so that the initial and final steps are east and north, respectively, and the height of the part of the lattice path within the strip $j-1\leq x\leq j$ is $a_{1}+a_{2}+\cdots+a_{j}$ for every $j\in[n]$. We observe that the number of nonzero entries of such an element of ${\mathbb{N}}^{n}$ is equal to one less than the number of east-north corners of the associated lattice path.
We denote by ${\mathcal{G}}_{n}$ the set of lattice paths which represent the elements of ${\mathbb{N}}^{n}$ with sum of coordinates at most $n$ and let $\Gamma_{\sigma}\in{\mathcal{G}}_{n}$ be the lattice path associated to the $n$-tuple with entry $s_{i}$ in position $s_{i-1}+1$ for $1\leq i\leq k$ (where $s_{0}:=0$) and all other entries equal to zero. Thus, $\Gamma_{\sigma}$ has successive steps east, followed by $r_{1}$ north and $r_{1}$ east steps, followed by $r_{2}$ north and $r_{2}$ east steps, and so on, followed by a string of north steps at the end. Then, the elements of $P_{\sigma}$ are represented by the lattice paths in ${\mathcal{G}}_{n}$ which lie weakly below $\Gamma_{\sigma}$.
@@
Let $\sigma=(1,1,\dots,1)\models n$, so that $P_{\sigma}$ consists of all points $(a_{1},a_{2},\dots,a_{n})\in{\mathbb{N}}^{n}$ such that $a_{1}+a_{2}+\cdots+a_{j}\leq j$ for every $j\in[n]$. The lattice path $\Gamma_{\sigma}$ has successive steps $\textrm{ENEN}\cdots\textrm{EN}$ and the total number of lattice paths in ${\mathcal{G}}_{n}$ which lie weakly below it (and hence the cardinality of $P_{\sigma}$) is well known to equal the Catalan number $\frac{1}{n+2}{2n+2\choose n+1}$. Moreover, counting these lattice paths by the number of east-north corners shows that $h_{i}(\sigma)$ is equal to the Narayana number
@@
(see, for instance, again [1, Theorem 2.32]). ∎
@@
([5, Conjecture 3.2]) We have $h(\sigma)=h({\rm rev}(\sigma))$ for every $\sigma\models n$.
@@
We observe that orthogonal reflection in the line $x+y=n$ maps $\Gamma_{\sigma}$ to $\Gamma_{{\rm rev}(\sigma)}$ and hence induces a bijection from the set of lattice paths which represent the elements of $P_{\sigma}$ to the set of those for $P_{{\rm rev}(\sigma)}$. We also observe that the number of east-north corners of a lattice path is preserved under this bijection and the proof follows. ∎

### v3

@@ Abstract.
An $n$-dimensional lattice polytope ${\mathcal{Q}}_{\sigma}$ can be associated to any composition $\sigma$ of a positive integer $n$, as a special case of constructions due to Pitman–Stanley and Chapoton. The entries of the $h$-vector of $\sigma$, introduced by Chapoton, enumerate the lattice points in ${\mathcal{Q}}_{\sigma}$ by the number of their nonzero coordinates. Chapoton conjectured that this vector is equal to the $h$-vector of a flag simplicial polytope. This paper proves this conjecture. Moreover, it shows that the gamma-vector associated to the $h$-vector of $\sigma$ is nonnegative by means of an explicit combinatorial interpretation and confirms certain other conjectures of Chapoton on the lattice point enumeration of composition polytopes. A combinatorial interpretation of their $h^{\ast}$-polynomials is deduced.
@@ 1. Introduction
The face lattice of composition polytopes is well understood. As observed in [16, Theorem 20], ${\mathcal{Q}}_{\sigma}$ is combinatorially isomorphic to a product of simplices, the dimensions of which are the parts of $\sigma$. The lattice point enumeration of composition polytopes, which is more interesting, has been studied for the more general classes of arbor polytopes [7] and Pitman–Stanley polytopes [16]. Chapoton defined the $h$-vector of a composition $\sigma$ of $n$ (more generally, of an arbor) as the sequence $h(\sigma)=(h_{i}(\sigma))_{0\leq i\leq n}$, where $h_{i}(\sigma)$ is the number of points in ${\mathcal{Q}}_{\sigma}\cap{\mathbb{N}}^{n}$ having exactly $i$ nonzero coordinates (this vector should not be confused with the $h$-vector of the simple polytope ${\mathcal{Q}}_{\sigma}$). We note that $h_{0}(\sigma)=h_{n}(\sigma)=1$. The following statement is a special case of an intriguing conjecture of Chapoton [7, Conjecture 0.1], stated for all arbor polytopes; we refer to the discussion in [7, Section 0.1] for the motivation behind it.
@@
In particular, $h(\sigma)$ should be palindromic and unimodal, meaning that $h_{i}(\sigma)=h_{n-i}(\sigma)$ for all $0\leq i\leq n$ and $h_{0}(\sigma)\leq h_{1}(\sigma)\leq\cdots\leq h_{\lfloor n/2\rfloor}(\sigma)$. Moreover, it is suggested in [7, Section 0.1] that $h(\sigma)$ should in fact be equal to the $h$-vector of an $n$-dimensional flag simplicial polytope (a simplicial polytope $\mathcal{Q}$} is said to be flag if every clique in the one-skeleton of $\mathcal{Q}$} is the vertex set of a face of $\mathcal{Q}$}). Thus, in view of Gal’s conjecture [9, Conjecture 2.1.7] (see also [2, Section 3]) on gamma-positivity, it is reasonable to expect that the generating polynomial $h(\sigma,t):=\sum_{i=0}^{n}h_{i}(\sigma)t^{i}$ of $h(\sigma)$ is $\gamma$-positive, meaning that
@@
for some nonnegative integers $\gamma_{i}(\sigma)$. The first main result of this paper states that all these claims are true for composition polytopes.
@@
The vector $h(\sigma)$ is equal to the $h$-vector of an $n$-dimensional flag simplicial polytope for every composition $\sigma$ of $n$. Moreover, $h(\sigma,t)$ is $\gamma$-positive and, in particular, palindromic and unimodal.
Our second main result confirms another curious conjecture of Chapoton [7], concerning the Ehrhart polynomial of composition polytopes. We recall (see, for instance, [14, Section 4.6]) that the Ehrhart polynomial ${\rm Ehr}({\mathcal{Q}},t)$ of a lattice polytope ${\mathcal{Q}}\subset{\mathbb{R}}_{\geq 0}^{n}$ has the defining property that ${\rm Ehr}({\mathcal{Q}},m)=\#(m{\mathcal{Q}}\cap{\mathbb{N}}^{n})$ for every $m\in{\mathbb{N}}$. The zeta polynomial (see [14, Section 3.12]) ${\mathcal{Z}}(P,t)$ of a finite poset $(P,\preceq)$ has the defining property that ${\mathcal{Z}}(P,m+1)$ is equal to the number of $m$-element multichains in $P$, meaning sequences $(p_{1},p_{2},\dots,p_{m})$ of elements of $P$ such that $p_{1}\preceq p_{2}\preceq\cdots\preceq p_{m}$, for every $m\in{\mathbb{N}}$. Given a composition $\sigma$ of $n$, Chapoton [7, Section 1] has considered the set $P_{\sigma}$ of lattice points in ${\mathcal{Q}}_{\sigma}$ as a subposet of the product poset ${\mathbb{N}}^{n}$.
@@
([7, Conjecture 3.1]) We have ${\rm Ehr}({\mathcal{Q}}_{\sigma},t)={\mathcal{Z}}(P_{{\rm rev}(\sigma)},t+1)$ for every composition $\sigma$, where ${\rm rev}(\sigma)$ is the composition obtained from $\sigma$ by reversing the order in which the parts of $\sigma$ appear.
Here is another intriguing conjecture from [7], stated there for all arbor polytopes (see [7, Conjecture 2.2]), on which no progress is made in this paper.
@@
Theorem 1.2 is proven in Section 3. This section shows that $h(\sigma,t)$ is equal to the Eulerian polynomial of a graded poset of width two (see Theorem 3.1) and deduces Theorem 1.2 from results of Brändén [6] and Reiner–Welker [11] on poset Eulerian polynomials. Section 2 provides a direct proof of the gamma-positivity of $h(\sigma,t)$ via an explicit combinatorial interpretation of the integers $\gamma_{i}(\sigma)$ which appear in Equation (2) (see Proposition 2.2); it also shows that $h(\sigma)=h({\rm rev}(\sigma))$ for every composition $\sigma$ (see Proposition 2.5). Two proofs of Theorem 1.3, one essentially bijective and another based on the results of [16, Section 3], are given in Section 4 and a combinatorial interpretation of the $h^{\ast}$-polynomial of ${\mathcal{Q}}_{\sigma}$ is deduced via the theory of pure shellability of posets (see Proposition 4.2). While this paper is restricted to composition polytopes, we do hope that some of the arguments presented here can be extended to more general classes of lattice polytopes.
Throughout this paper, we set ${\mathbb{N}}=\{0,1,2,\dots\}$ and $n]=\{1,2,\dots,n\}$ for $n\in{\mathbb{N}}$. We denote by #S\#S the cardinality of a finite set $S$ and write $\sigma\models n$ when $\sigma$ is a composition of $n$. Our notation and terminology about posets follows [14, Chapter 3]. Background on Ehrhart theory and face enumeration of simplicial complexes and polytopes can be found in [4, 10, 13] [14, Chapter 4].
@@ 2. Gamma-positivity
This section provides a direct combinatorial proof of the gamma-positivity of $h(\sigma,t)$ and shows that $h(\sigma)=h({\rm rev}(\sigma))$ for every composition $\sigma\models n$. Moreover, it introduces a lattice path interpretation of points in $P_{\sigma}={\mathcal{Q}}_{\sigma}\cap{\mathbb{N}}^{n}$ which will be useful in the following section as well.
@@
(see [2, Theorem 2.32] for a generalization in the setting of finite Coxeter groups). ∎
Lattice path interpretation. For the remainder of this section we represent elements of ${\mathbb{N}}^{n}$ with sum of coordinates at most $n$ by lattice paths in the $xy$-plane as follows. A point $(a_{1},a_{2},\dots,a_{n})\in{\mathbb{N}}^{n}$ with $a_{1}+a_{2}+\cdots+a_{n}\leq n$ is represented by the lattice path with initial vertex $(0,0)$, terminal vertex $(n,n)$ and east (E) or north (N) unit steps, so that the height of the part of the lattice path within the strip $j-1\leq x\leq j$ is $a_{1}+a_{2}+\cdots+a_{j}$ for every $j\in[n]$. We observe that the number of nonzero entries of such an element of ${\mathbb{N}}^{n}$ is equal to the number of north-east corners of the associated lattice path.
We denote by ${\mathcal{G}}_{n}$ the set of lattice paths which represent the elements of ${\mathbb{N}}^{n}$ with sum of coordinates at most $n$ and let $\Gamma_{\sigma}\in{\mathcal{G}}_{n}$ be the lattice path associated to the $n$-tuple with entry $s_{i}$ in position $s_{i-1}+1$ for $1\leq i\leq k$ (where $s_{0}:=0$) and all other entries equal to zero. Thus, $\Gamma_{\sigma}$ has $r_{1}$ north steps, followed by $r_{1}$ east steps, followed by $r_{2}$ north steps, followed by $r_{2}$ east steps, and so on. Then, the elements of $P_{\sigma}$ are represented by the lattice paths in ${\mathcal{G}}_{n}$ which lie weakly below $\Gamma_{\sigma}$.
@@
Let $\sigma=(1,1,\dots,1)\models n$, so that $P_{\sigma}$ consists of all points $(a_{1},a_{2},\dots,a_{n})\in{\mathbb{N}}^{n}$ such that $a_{1}+a_{2}+\cdots+a_{j}\leq j$ for every $j\in[n]$. The lattice path $\Gamma_{\sigma}$ has successive steps $\textrm{NENE}\cdots\textrm{NE}$ and the total number of lattice paths in ${\mathcal{G}}_{n}$ which lie weakly below it (and hence the cardinality of $P_{\sigma}$) is well known to equal the Catalan number $\frac{1}{n+2}{2n+2\choose n+1}$. Moreover, counting these lattice paths by the number of north-east corners shows that $h_{i}(\sigma)$ is equal to the Narayana number
@@
(see, for instance, again [2, Theorem 2.32]). ∎
@@
([7, Conjecture 3.2]) We have $h(\sigma)=h({\rm rev}(\sigma))$ for every $\sigma\models n$.
@@
We observe that orthogonal reflection in the line $x+y=n$ maps $\Gamma_{\sigma}$ to $\Gamma_{{\rm rev}(\sigma)}$ and hence induces a bijection from the set of lattice paths which represent the elements of $P_{\sigma}$ to the set of those for $P_{{\rm rev}(\sigma)}$. We also observe that the number of north-east corners of a lattice path is preserved under this bijection and the proof follows. ∎
3. Polytopality
This section proves Theorem 1.2. We first need to recall some background and terminology about the face enumeration of simplicial polytopes and the combinatorics of posets; more information can be found in [11, 13] [14, Section 3.15].
The $h$-polynomial of a $d$-dimensional simplicial polytope $\mathcal{Q}$} is a convenient way to record the face numbers of $\mathcal{Q}$}; it is defined by the formula
(6) $h_{\mathcal{Q}}(t)=\sum_{i=0}^{d}f_{i-1}({\mathcal{Q}})t^{i}(1-t)^{d-i},$
where $f_{i-1}({\mathcal{Q}})$ is the number of (i1)(i-1)-dimensional faces of $\mathcal{Q}$}. If $h_{\mathcal{Q}}(t)=\sum_{i=0}^{d}h_{i}({\mathcal{Q}})t^{i}$, then $(h_{0}({\mathcal{Q}}),h_{1}({\mathcal{Q}}),\dots,h_{d}({\mathcal{Q}}))$ is called the $h$-vector of $\mathcal{Q}$}. The polynomial $h_{\mathcal{Q}}(t)$ has nonnegative coefficients and, as a consequence of the $g$-theorem [13, Section III.1], it is palindromic and unimodal for every simplicial polytope $\mathcal{Q}$}. Moreover, it is conjectured to be $\gamma$-positive [9, Conjecture 2.1.7] whenever $\mathcal{Q}$} is flag.
Let $P$ be a poset with $p$ elements, with partial order denoted by $\preceq$, and $\omega:P\rightarrow[p]$ be a natural labeling, meaning a bijection with the property that $a\prec b\Rightarrow\omega(a)<\omega(b)$ for all $a,b\in P$. A linear extension of $P$ is defined as any permutation $(a_{1},a_{2},\dots,a_{p})$ of $P$ such that $a_{i}\prec a_{j}\Rightarrow i<j$ for all $i,j\in[p]$. The Jordan-Hölder set ${\mathcal{L}}(P,\omega)$ consists of all permutations $(\omega(a_{1}),\omega(a_{2}),\dots,\omega(a_{p}))$ of $[p]$, where $(a_{1},a_{2},\dots,a_{p})$ is a linear extension of $P$. The Eulerian polynomial of $P$ is defined as
(7) $W_{P}(t)=\sum_{w\in{\mathcal{L}}(P,\omega)}t^{{\rm des}(w)},$
where ${\rm des}(w)$ stands for the number of indices $i\in[p-1]$ (called descents) such that $w_{i}>w_{i+1}$, if $w=(w_{1},w_{2},\dots,w_{p})$. This polynomial is independent of the natural labeling $\omega$; it was shown to be $\gamma$-positive and equal to the $h$-polynomial of a (pd)(p-d)-dimensional simplicial polytope by Brändén [6] and Reiner–Welker [11], respectively, for every graded poset $P$, where $d$ is the cardinality of any maximal chain of $P$.
The width of a poset $P$ is defined as the largest size of an antichain (subset of pairwise incomparable elements) in $P$. The following statement is a stronger version of Theorem 1.2.
Theorem 3.1.
The polynomial $h(\sigma,t)$ is equal to the Eulerian polynomial of a graded poset of width two for every composition $\sigma\models n$.
In particular, $h(\sigma,t)$ is $\gamma$-positive, hence palindromic and unimodal, and equal to the $h$-polynomial of an $n$-dimensional flag simplicial polytope.
Proof.
Let $\sigma=(r_{1},r_{2},\dots,r_{k})\models n$ and $s_{i}=r_{1}+r_{2}+\cdots+r_{i}$ for $i\in[k]$ be as in Section 2. We consider a poset $\Pi_{\sigma}$ which consists of two disjoint chains $\alpha_{1}\prec\alpha_{2}\prec\cdots\prec\alpha_{n}$ and $\beta_{1}\prec\beta_{2}\prec\cdots\prec\beta_{n}$, together with the additional cover relations $\alpha_{s_{i}}\prec\beta_{s_{i}+1}$ for every $i\in[k-1]$. This poset is graded; all maximal chains have cardinality $n$. We also consider the natural labeling $\omega:\Pi_{\sigma}\rightarrow[2n]$ such that $\omega(\alpha_{i})=i$ and $\omega(\beta_{i})=n+i$ for $i\in[n]$.
We claim that $h(\sigma,t)=W_{\Pi_{\sigma}}(t)$. Indeed, we recall from Section 2 that the coefficient $h_{j}(\sigma)$ of $t^{j}$ in $h(\sigma,t)$ is equal to the number of lattice paths in ${\mathcal{G}}_{n}$ which lie weakly below $\Gamma_{\sigma}$ and have exactly $j$ north-east corners. On the other hand, the coefficient of $t^{j}$ in $W_{\Pi_{\sigma}}(t)$ is equal to the number of permutations $w\in{\mathcal{L}}(\Pi_{\sigma},\omega)$ such that ${\rm des}(w)=j$. To any linear extension of $\Pi_{\sigma}$ one can associate the lattice path in ${\mathcal{G}}_{n}$ obtained by replacing each element $\alpha_{i}$ by an east step and each element $\beta_{i}$ by a north step. These lattice paths lie weakly below $\Gamma_{\sigma}$ and the resulting map induces a one-to-one correspondence between permutations $w\in{\mathcal{L}}(\Pi_{\sigma},\omega)$ and lattice paths in ${\mathcal{G}}_{n}$ which lie weakly below $\Gamma_{\sigma}$. Moreover, the number of descents of $w\in{\mathcal{L}}(\Pi_{\sigma},\omega)$ is equal to the number of north-east corners of the corresponding lattice path. These facts show that $t^{j}]W_{\Pi_{\sigma}}(t)=h_{j}(\sigma)$ for every $j$ and the proof of the claim follows. The first statement in the theorem follows from the claim. The second statement follows from the first and the aforementioned results of [6, 11]. Since $\Pi_{\sigma}$ has width two, the simplicial polytope with $h$-polynomial $h(\sigma,t)$ can be chosen to be flag by [11, Theorem 3.23]. ∎
Remark 3.2.
After this paper first appeared on the arXiv, Frederic Chapoton [8] suggested to the author that $h(\sigma,t)$ may be equal to the $h^{\ast}$-polynomial of the order polytope of the poset $\Pi_{\sigma}$. This suggestion led to Theorem 3.1.
The lattice path interpretation of $h(\sigma,t)$ shows that it belongs to the class of nonnesting rook polynomials, introduced and studied recently by Alexandersson and Jal [1]. Given that, Theorem 3.1 can also be deduced from the results of [1, Section 5]. Corollary 4.2 in [1] implies that $h(\sigma,t)$ has another important property, namely that it is ultra log-concave for every composition $\sigma$. To the best of our knowledge, the real-rootedness of $h(\sigma,t)$ is open.